OpenStudy (anonymous):

A square garden has a side of (x+7y)m.What would be the effect on its area if the length of each side is doubled? halved? explain

OpenStudy (anonymous):

Let's simplify things by saying that the side length is \(s\). Now, the area is \(A = s^2\). If instead we had \(2s\) (double the original value), the new area would be \(A'=(2s)^2=4s^2=4A\), and if we had \(s/2\) (half the original value) the new area would be \(A''=(s/2)^2=s^2/2=A/4\).

OpenStudy (anonymous):

Typo fix: \(A''=(s/2)^2=s^2/\color{red} 4=A/4\)